OpenStudy (anonymous):

If log16(4) = x what is x A) 4 B) 3 c) 2 D ) 1 E) 1.5

OpenStudy (anonymous):

any of this answers are correct :o

OpenStudy (anonymous):

correct answer is 1/2 = 0.5

OpenStudy (anonymous):

ohh i thouhg u say all the answer correct?

OpenStudy (anonymous):

if you want to do this equation in your calculator just do log(4)/log(16) all in base 10

OpenStudy (anonymous):

ohh okay thank you very much

OpenStudy (anonymous):

\[\log _{16}x=4\]\[16^x=4\]\[2^{4\times x } =2^2\]\[4x=2\]\[x=\frac{2}{4}\]\[x=\frac{1}{2}\]